OpenStudy (mackenzie2013):

There are 3.76 × 1022 atoms in 1 gram of oxygen. How many atoms are there in 700 grams of oxygen? Write your answer in scientific notation. If necessary, round your answer to two decimal places.

OpenStudy (zale101):

Have you started doing the problem? What have you got so far?

OpenStudy (zale101):

Hint: Yo have a conversion factor already been given to you. "There are 3.76 × 1022 atoms in 1 gram of oxygen. \(\LARGE\frac{3.76*10^{22} ~atoms}{1~gram~Oxygen}\) or \(\LARGE\frac{1~gram~Oxygen}{3.76*10^{22} ~atoms}\)

OpenStudy (mackenzie2013):

I dont understand it all around.

OpenStudy (campbell_st):

its a simple multiplication \[(3.76 \times 10^{22}) \times 700\] if your calculator is in scientific mode it should be able to handle this